OpenStudy (anonymous):

Solve the equation for y. 4x – 5y = 9 A.y=-5/4x-5/9 B.y=5/4x+9/4 C.y=4/5x-9/5 D.y=-4x-9

OpenStudy (igreen):

4x – 5y = 9 Subtract 4x to both sides: -5y = -4x + 9 Can you divide -5 to all terms? @NunChuckBoy3101

OpenStudy (anonymous):

i think it's c

OpenStudy (anonymous):

i think so im not great at math

OpenStudy (igreen):

Yep, you got it!

OpenStudy (anonymous):

so i am right it is c

OpenStudy (igreen):

Dividing -5 to all terms will give you: \(y = \dfrac{4}{5}x - \dfrac{9}{5}\)

OpenStudy (igreen):

Yes, you're right.

OpenStudy (anonymous):

thank you will you help me with another

OpenStudy (igreen):

Sure, just tag me.